OpenStudy (anonymous):

8-20). Simplify the following radical expression. (Enter your answer as an improper fraction reduced to the lowest terms.) root3(27/8)

OpenStudy (anonymous):

cbrt(27/8)=3/2=1.5

OpenStudy (anonymous):

\[\sqrt[3]{27/8}\] = 3/2

OpenStudy (anonymous):

take the cubed root of the numerator and the cubed root of the denominator